40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1572     Entry time: Sun May 10 13:41:17 2009
Author: steve 
Type: Update 
Category: VAC 
Subject: ETMY damping restored, VC1 opened 

ETMY damping restored.

Cryo  interlock closed VC1 ~2 days ago. P1 is 6.3 mTorr. Cryo temp 12K stable, reset photoswitch and opened VC1

ELOG V3.1.3-